a woman has $2.65 in dimes & nickels. she has 13 more dimes than nickels. how many nickels does she have

Answer by stanbon(75887)

a woman has $2.65 in dimes & nickels. she has 13 more dimes than nickels. how many nickels does she have
-------
Equations:
Quantity Eq:: d - n = 13 coins
Value Eq:: 10d + 5n = 265 cents
-------------------------
Modify for elimination:
d - n = 13
2d + n = 53
-------------------
Add and solve for "d":
3d = 66
d = 22 (# of dimes)
----
Solve for "n"
n = d-13
n = 9 (# of nickels)
---------------------------
Cheers,
Stan H.
---------------